OpenStudy (anonymous):

The ratio of milk and water in a mixture is 2 : 3 and the ratio of water to milk in another mixture is 3 : 4. If 70 litres of each type of milk is added to 100 litres of pure milk, what will be the ratio of milk and water now?

OpenStudy (perl):

you can make a proportion

OpenStudy (anonymous):

2*70/5=28,3*70/5=42,3*70/7=30,4*70/7=40 so ratio of milk is 58 in 70 litres and water is 82added to pure milk is 100+58=158 milk,82 water ratio 158:82

OpenStudy (anonymous):

is it right?

OpenStudy (callisto):

<Type A>The ratio of milk to water = 2 : 3 <Type B>The ratio of water to milk = 3 : 4. <I> We have 70L of <Type A>: Milk = 70 * (2/5) = 28 Water = 70 - 28 = 42 <II> We have 70L of <Type B> Milk = 70 * (3/7) = 30 Water = 70 - 30 = 40 Total milk = 28 + 30 = 58 Total water = 42 + 40 = 82 Adding 100L milk... Total milk = 58 + 100 = 158 Total water = unchanged = 82 Ratio of milk to water = 158 : 82 Good, but that's not the end. You need to simplify 158 : 82!

OpenStudy (anonymous):

how to simplify 79:41 ?

OpenStudy (callisto):

79 : 41 is right :)

OpenStudy (callisto):

WAIIITTTT

OpenStudy (anonymous):

no i am not getting the answer its in simple ratio like 3:4

OpenStudy (callisto):

<Type A>The ratio of milk to water = 2 : 3 <Type B>The ratio of water to milk = 3 : 4. <I> We have 70L of <Type A>: Milk = 70 * (2/5) = 28 Water = 70 - 28 = 42 <Corrected> <II> We have 70L of <Type B> Milk = 70 * (4/7) = 40 Water = 70 - 40 = 30 Total milk = 28 + 40 = 68 Total water = 42 + 30 = 72 Adding 100L milk... Total milk = __ + 100 = ... Total water = unchanged = 72 Ratio of milk to water = ...?

OpenStudy (callisto):

Type B is the ratio of water to milk = 3 : 4. I was sooooo careless!! :(

OpenStudy (anonymous):

hey i also did that wrong right it is 7:3.i got it.thank u very much

OpenStudy (callisto):

You're welcome! We have to be careful next time!

OpenStudy (anonymous):

yesterdY I saw that thing but today i forget that it is water to milk .so for confusing they gave this type of questions

OpenStudy (callisto):

Practice will help. I never knew how to do this type of question when I was taught about this. :\

OpenStudy (anonymous):

i have another question

OpenStudy (callisto):

Close this question and ask in a new post :)

OpenStudy (anonymous):

k